# Symmetric Preparation of Systems

### v2

@@
Abstract.
In this paper we generalize the Weierstrass and Malgrange preparation theorems to the symmetric matrix valued case, proving symmetric preparation of analytic and smooth symmetric systems that vanishes of first order.
2010 Mathematics Subject Classification:
Primary 32B05 Secondary 47A56, 26B40
@@ 1. Introduction and statement of results
If $f$ is $C^{\infty}$ satisfying $\partial_{t}f(0,0)\neq 0$ and $f(0,0)=0$ then by the Malgrange preparation theorem there exist $C^{\infty}$ functions $c(t,x)$ and $r(x)$ such that $c(0,0)\neq 0$, $r(0)=0$ and
(1.1) $f(t,x)=c(t,x)\left(t+r(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$. In the analytic case, the Weierstrass preparation theorem, one has uniqueness but not in the $C^{\infty}$ case, except for real valued $f(x)$ when one can use the implicit function theorem. This also holds for higher order vanishing of $f$ and there are corresponding division results.
The Malgrange preparation theorem has been generalized to $N\times N$ systems of functions in [3]. The simplest example is for smooth $N\times N$ system $F(t,x)$ satisfying $|\partial_{t}F(0,0)|\neq 0$ and $F(0,0)=0$. Then there exist smooth $N\times N$ systems $C(t,x)$ and $M(x)$ such that $|C(0,0)|\neq 0$, $M(0)=0$ and
(1.2) $F(t,x)=C(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$, see [2]. In the analytic case, one has uniqueness but not in the smooth case.
We also proved the following division theorem in [2]. If $F(t,x)$ is a smooth $N\times N$ system, then for any smooth $N\times N$ system $M(x)$ there exist smooth $N\times N$ systems $C(t,x)$ and $R(x)$ such that
(1.3) $F(t,x)=C(t,x)\left(t\operatorname{Id}+M(x)\right)+R(x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$. As before, one has uniqueness in the analytic case but not in the smooth case. We also have preparation results for higher order vanishing of $F$, adjoint preparation results and corresponding division results, see [3].
Now, the question about symmetric preparation results was raised by Gregory Berkolaiko [1]. The simplest case is for smooth (or analytic) symmetric $N\times N$ matrices $F(t,x)$ satisfying $\partial_{t}F(0,0)>0$ and $F(0,0)=0$. Then, does it exist smooth (or analytic) $N\times N$ matrices $U(t,x)$ and $M(x)$, such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric with $M(0)=0$ satisfying
@@
near $(0,0)$, and is this preparation unique? Observe that the conditions on $F(t,x)$ are necessary for (1.4) to hold since $U(0,0)U^{*}(0,0)=\partial_{t}F(0,0)>0$ and $F(0,0)=U(0,0)M(0)U(0,0)^{*}=0$. Here and in the following we say that the matrix $A$ is symmetric if $A^{*}=A$ and antisymmetric if $A^{*}=-A$, where $A^{*}=\overline{A}^{t}$.
@@
Observe that the preparation (1.4) is not unique in general. In fact, if $U(t,x)$ and M(x) are solutions to (1.4), then for any orthogonal $A(x)=(A^{-1})^{*}(x)$ we find that $\widetilde{U}(t,x)=U(t,x)A(x)$ and $\widetilde{M}(x)=A^{*}(x)M(x)A(x)$ solves (1.4).
Actually, the preparation is unique in the case when $U$ is symmetric and $U>0$, see Theorem 1.2. We are going to prove the following, which is the main result of the paper.
@@
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ such that
(1.5) $\partial_{t}F(0,0)>0\quad\text{and}\quad F(0,0)=0$
Then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M(x)$ such that $|U|\neq 0$, $M$ is symmetric , $M(0)=0$ and
(1.6) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)\qquad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.6) with analytic $U(t,x)$ and $M(x)$. This preparation is unique if $U$ is symmetric and $U>0$.
Remark 1.3.
Observe that we also get uniqueness in other cases, for example if $U<0$ by taking U-U for for the unique preparations with $U>0$. The proof of Theorem 1.2 also gives the preparation (1.6) for some $0\neq|F(0,0)|\ll 1$ with $|M(0)|\ll 1$, but the bounds depend on the (semi)norms of $F$.
If $F(0,0)\neq 0$ then by subtracting $F(0,0)$ we obtain (1.6) with the remainder $F(0,0)$ giving the following result.
Corollary 1.4.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ such that
(1.7) $\partial_{t}F(0,0)>0$
Then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M(x)$ such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric $\forall\,x$, $M(0)=0$ and
(1.8) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)+F(0,0)\qquad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.8) with analytic $U(t,x)$ and $M(x)$. This preparation is unique if $U>0$.
Note that the uniqueness only holds for symmetric preparations, since Remark 1.1 gives counterexamples for nonsymmetric $U$. Since (1.6) is not linear in $U(t,x)$ the proof approaches of Weierstrass or Malgrange preparation results do not work in the symmetric case, instead we shall use inverse function theorems. Theorem 1.2 will be proved in Section 3 for the analytic case, and in Section 4 for the smooth case. But first we shall consider the formal power series case.
2. Formal power series solutions
If $F$ is analytic we may solve (1.4) by taking the Taylor expansion of $F(t,x)$:
(2.1) $F(t,x)=\sum_{{j,\alpha}}F_{j,\alpha}t^{j}x^{\alpha}$
where $F_{0,0}=F(0,0)=0$ and $F_{1,0}=\partial_{t}F(0,0)>0$, and take the formal power series expansions $U(t,x)=\sum_{{j,\alpha}}U_{\begin{subarray}{c}j,\alpha\end{subarray}}t^{j}x^{\alpha}$ where $|U_{0,0}|\neq 0$ and $M(x)=\sum_{\alpha\neq 0}M_{\alpha}x^{\alpha}$ is symmetric such that $M_{0}=M(0)=0$. Then (1.4) gives that
(2.2) $\sum_{{j,\alpha}}F_{j,\alpha}t^{j}x^{\alpha}=\sum_{{j,\alpha}}U_{j,\alpha}t^{j}x^{\alpha}\left(t\operatorname{Id}_{N}+\sum_{\gamma\neq 0}M_{\gamma}x^{\gamma}\right)\sum_{{k,\beta}}U_{k,\beta}^{*}t^{k}x^{\beta}$
Simplifying the notations so that $\sum_{{j,\alpha}}A_{\begin{subarray}{c}j,\alpha\end{subarray}}t^{j}x^{\alpha}=\sum_{j,k}A_{j,k}t^{j}x^{k}$, we get
(2.3) $\sum_{j,k}F_{j,k}t^{j}x^{k}=\sum_{j,k}U_{j,k}t^{j}x^{k}\left(t\operatorname{Id}_{N}+\sum_{m>0}M_{m}x^{m}\right)\sum_{{i,\ell}}U_{i,\ell}^{*}t^{i}x^{\ell}$
and by matching coefficients we obtain that
(2.4) $F_{j,k}=\sum_{\begin{subarray}{c}i<j\\ \ell\leq k\end{subarray}}U_{i,\ell}U_{j-1-i,k-\ell}^{*}+\sum_{\begin{subarray}{c}i\leq j\\ \ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{i,\ell}M_{m}U_{j-i,k-m-\ell}^{*}$
The first terms of this expansion are $F_{0,0}=0$, $F_{0,1}=U_{0,0}M_{1}U_{0,0}^{*}$ and $F_{1,0}=U_{0,0}U_{0,0}^{*}$ which gives that $|U_{0,0}|\neq 0$ and $M_{1}=U_{0,0}^{-1}F_{0,1}(U_{0,0}^{-1})^{*}$.
When $j=0$ we obtain from (2.4) that
(2.5) $F_{0,k}=\sum_{\begin{subarray}{c}\ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}\qquad\forall\,k$
so recursion gives
(2.6) $U_{0,0}M_{k}U_{0,0}^{*}=F_{0,k}-\sum_{\begin{subarray}{c}0\leq\ell\leq k-m\\ 0<m<k\end{subarray}}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}\qquad\forall\,k$
We then obtain symmetric $M_{j}$ by
(2.7) $M_{k}=U_{0,0}^{-1}F_{0,k}(U_{0,0}^{-1})^{*}-\sum_{\begin{subarray}{c}0\leq\ell\leq k-m\\ 0<m<k\end{subarray}}U_{0,0}^{-1}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}(U_{0,0}^{-1})^{*}\qquad\forall\,k$
so $M_{k}$ is determined by $F_{0,k}$, $U_{0,\ell}$ with $\ell<k$ and $M_{m}$ with $m<k$.
In general, we obtain from (2.4) with $j$ replaced by $j+1$ that
(2.8) $2\operatorname{Re}U_{j,k}U_{0,0}^{*}=F_{j+1,k}-\sum_{\begin{subarray}{c}0<i<j\\ \ell\leq k\end{subarray}}U_{i,\ell}U_{j-1-i,k-\ell}^{*}-\sum_{\begin{subarray}{c}i\leq j+1\\ \ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{i,\ell}M_{m}U_{j+1-i,k-m-\ell}^{*}$
where the right hand side is determined by $F_{j+1,k}$, $M_{m}$ with $m\leq k$ determined by (2.6) and $U_{i,\ell}$ with $i<j$ and $\ell\leq k$ or $i\leq j+1$ and $\ell\leq k-1$, which gives $i+\ell<j+k$ and $\ell\leq k$ or $i+\ell\leq j+k$ and $\ell\leq k-1$.
Since $|U_{0,0}|\neq 0$, (2.8) can be solved to give $U_{j,k}$ and by Lemma A.1 we get a unique symmetric solution $U_{j,k}$ when $U_{0,0}$ is symmetric. For $k=0$ we find from (2.8) that
(2.9) $2\operatorname{Re}U_{j,0}U_{0,0}^{*}=F_{j+1,0}-\sum_{\begin{subarray}{c}0<i<j\end{subarray}}U_{i,0}U_{j-1-i,0}^{*}$
which gives $2\operatorname{Re}U_{1,0}U_{0,0}^{*}=F_{2,0}$. Since $|U_{0,0}|\neq 0$ we obtain by induction $U_{j,0}$ for any $j$ from (2.9), and if $U_{0,0}$ is symmetric then Lemma A.1 gives a unique symmetric $U_{j,0}$.
Thus, one can determine $U_{j,k}$ and $M_{k}$ from (2.7) and (2.8) by induction over $j+k=p\geq 0$. For $p=0$ we can take any $U_{0,0}$ such that $U_{0,0}U_{0,0}^{*}=F_{1,0}>0$. Then for fixed $p>0$ we get the values of $U_{p,0}$ from (2.9) so we can do a recursion over (pr,r)(p-r,r) for $r=0$ to $p$. If we have determined $U_{j,k}$ for $j+k<p$ and $U_{p-\ell,\ell}$ for $\ell\leq r$ for some $r<p$ then by (2.8) we find that $U_{p-r-1,r+1}$ only depends on $F_{p-r,r+1}$, $U_{i,\ell}$ for $i+\ell\leq p$ and $\ell\leq r$ or $i+\ell<p$ and $\ell\leq r+1$, which are previously determined, and on $M_{m}$ with $m\leq r+1$ which by (2.7) are given by $F_{0,m}$ and $U_{0,\ell}$ with $\ell\leq r<p$ which also have been determined. By recursion over $r$ and induction over $p$ we get the expansion, and if $U_{0,0}$ is symmetric we get unique symmetric $U_{j,k}$ by (2.8) and Lemma A.1, which gives the following result.
Theorem 2.1.
@@ 2. Formal power series solutions
If $F(t,x)$ is analytic and symmetric such that $F(0,0)>0$ and $F(0,0)=0$ then there exists formal power series $U(t,x)$ and $M(x)$ solving (1.4) modulo $O(|t|^{N}+|x|^{N})$, $\forall\,N$. We get a unique power series solution in the case when $U$ is symmetric.
@@
But is it not clear that the formal power expansion converge to an analytic function.
Remark 2.2.
If $n=1$ and $F$ is analytic, then by  [1] the nested Artin approximation (see e.g. Theorem 5.18 in [6]) gives analytic $\overline{U}(x,y)=\sum_{j,k}\overline{U}_{jk}t^{j}x^{k}$ and $\overline{M}(x)=\sum_{k}\overline{M}_{k}x^{k}=\overline{M}^{*}(x)$ that solves
$F(t,x)=\overline{U}(t,x)(tId+\overline{M}(x))\overline{U}^{*}(t,x)\qquad(t,x)\in\mathbf{R}^{2}$
which gives Theorem 1.2 in the two dimensional analytic case. One also obtains that any finite number of terms in the expansions of $\overline{U}(t,x)$ and $\overline{M}(x)$ coincide with the corresponding terms in formal power series solutions $U(t,x)$ and $M(x)$.
When $n>1$ this does not work in general, see the counterexamples in [6]. Instead we will use the inverse function theorem, see Section 3.
@@
It is clear that $N\times N$ analytic systems form a Banach space with the usual $L^{{}^{\infty}}$ norm, and they can be localized to neighborhoods. We shall prove that if $U(t,x)$ and $M(x)$ are analytic such that $|U|\neq 0$, $M$ is symmetric and $M(0)=0$, then the mapping
(3.1) $(U(t,x),M(x))\overset{\mathcal{F}}{\longrightarrow}\Big(\partial_{t}\big(U(t,x)(t\operatorname{Id}_{N}+M(x))U^{*}(t,x)\big),U_{0}(x)M(x)U_{0}^{*}(x)\Big)\\ =(F_{1}(t,x),F_{0}(x))$
@@
with $U_{0}(x)=U({0,x})$, is analytic and surjective in some neighborhood of $(0,0)$ for analytic and symmetric $F_{j}$, j=0, 1j=0,\,1, when $|F_{1}(0,0)|\neq 0$ and $F_{0}(0)=0$. If we obtain a solution to (3.1) in a neighborhood of $(0,0)$ then we obtain (1.6) with
(3.2) $F(t,x)=F_{0}(x)+\int_{0}^{t}F_{1}(s,x)\,ds\qquad\text{near $(0,0)$}$
thus $F_{0}(x)=F(0,x)$ and $F_{1}(t,x)=\partial_{t}F(t,x)$.
When $M\equiv 0$ we find that $\partial_{t}(tUU^{*})=F_{1}$ so that
(3.3) $U(t,x)U^{*}(t,x)=\frac{1}{t}\int_{0}^{t}F_{1}(s,x)\,ds\to F_{1}(0,x)\text{ when }t\to 0$
@@
For any such solution $U(t,x)$, we find that the differential of $\mathcal{F}$ at $(U,0)$ is given by
(3.4) $(u(t,x),m(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}\big(2t\operatorname{Re}U(t,x)u^{*}(t,x)+U(t,x)m(x)U^{*}(t,x)\big),U_{0}(x)m(x)U_{0}^{*}(x)\Big)\\ =(A_{1}(t,x),A_{0}(x))$

### v3

# arXiv:2601.12458v3[math.CA] 28 Feb 2026

@@
arXiv:2601.12458v3[math.CA] 28 Feb 2026
Symmetric Preparation of Systems
Nils Dencker
Lund University
Department of Mathematics, University of Lund, Box 118, S-221 00 Lund, Sweden nils.dencker@gmail.com
@@ 1. Introduction and statement of results
If $f$ is $C^{\infty}$ satisfying $\partial_{t}^{m}f(0,0)\neq 0$ and $\partial_{t}^{j}f(0,0)=0$ for $j<m$, then by the Malgrange preparation theorem there exist $C^{\infty}$ functions $c(t,x)$ and $r(t,x)=\sum_{j=0}^{m-1}t^{j}r_{j}(x)$ such that $c(0,0)\neq 0$, $r_{j}(0)=0,\ \forall\,j$ and
(1.1) $f(t,x)=c(t,x)\left(t^{m}+r(t,x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$. In the analytic case, the Weierstrass preparation theorem, one has uniqueness but not in the $C^{\infty}$ case, except for $m=1$ and real valued $r(t,x)$ when one can use the implicit function theorem. There are also corresponding division results.
The Malgrange preparation theorem has been generalized to $N\times N$ systems of functions in [3]. An example is for smooth $N\times N$ system $F(t,x)$ satisfying
(1.2) $|\partial_{t}^{m}F(0,0)|>0\quad\text{and}\quad\partial_{t}^{j}F(0,0)=0,\quad j<m$
Then there exist smooth $N\times N$ systems $C(t,x)$ and $M_{j}(x)$, $j<m$, such that $|C(0,0)|\neq 0$, $M_{j}(0)=0$, $\forall\,j$, and
(1.3) $F(t,x)=C(t,x)\left(t^{m}\operatorname{Id}_{N}+\sum_{j=0}^{m-1}t^{j}M_{j}(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$, this is the special case of [3, Th. 5.3]. In the analytic case, one has uniqueness but not in the smooth case.
We also proved the following division theorem in [3]. If $F(t,x)$ is a smooth $N\times N$ system satisfying (1.2), then for any smooth $N\times N$ system $G(t,x)$ there exist smooth $N\times N$ systems $C(t,x)$ and $R_{j}(x)$, $j<m$, such that
(1.4) $G(t,x)=C(t,x)F(t,x)+\sum_{j=0}^{m-1}t^{j}R_{j}(x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$, this is a special case of [3, Th. 5.9]. As before, one has uniqueness in the analytic case but not in the smooth case. By taking the adjoints, we also obtain adjoint preparation results and division results, see [3].
Now, the question about symmetric preparation results was raised by Gregory Berkolaiko [1]. The simplest case is for smooth (or analytic) symmetric $N\times N$ matrices $F(t,x)$ satisfying $\partial_{t}F(0,0)>0$ and $F(0,0)=0$. Then the question was: does it exist smooth (or analytic) $N\times N$ matrices $U(t,x)$ and $M(x)$, such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric with $M(0)=0$, satisfying
@@
near $(0,0)$, and is this preparation unique? Observe that the conditions on $F(t,x)$ are necessary for (1.5) to hold since $U(0,0)U^{*}(0,0)=\partial_{t}F(0,0)>0$ and $F(0,0)=U(0,0)M(0)U^{*}(0,0)$. Here and in the following we say that the matrix $A$ is symmetric if $A^{*}=A$ and antisymmetric if $A^{*}=-A$, where $A^{*}=\overline{A}^{t}$, and when we say that $A>0$ then we assume that $A$ is symmetric. This result was proved in [4], and in this paper we shall generalize to the higher order case. The following is the main result of the paper.
Theorem 1.1.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ satisfying
(1.6) $\partial_{t}^{m}F(0,0)>0\quad\text{and}\quad\partial_{t}^{j}F(0,0)=0,\quad j<m$
then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M_{j}(x)$, $j<m$, such that $|U(0,0)|\neq 0$, $M_{j}(x)$ is symmetric such that $M_{j}(0)=0,\ \forall\,j,$ and
(1.7) $F(t,x)=U(t,x)\left(t^{m}\operatorname{Id}_{N}+\sum_{j=0}^{m-1}t^{j}M_{j}(x)\right)U^{*}(t,x)\quad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.7) with analytic $U(t,x)$ and ${\mathbb{M}}(t,x)$. One can choose $U>0$ and then the preparation is unique.
@@
Observe that the preparation (1.7) is not unique in general. For example, if $U(t,x)$ and ${\mathbb{M}}(t,x)$ are solutions to (1.7), then for any orthogonal $A(x)=(A^{-1})^{*}(x)$ we find that $\widetilde{U}(t,x)=U(t,x)A(x)$ and $\widetilde{\mathbb{M}}(t,x)=A^{*}(x){\mathbb{M}}(t,x)A(x)$ solves (1.7). But if $U(t,x)>0$ and $UA(t,x)>0$ then $A(t,x)\equiv\operatorname{Id}_{N}$.
In fact, if $U>0$ then $U$ has a ON base of eigenvectors, and if $UA$ is symmetric then $A^{*}U=UA$ which gives $U=AUA$. Then $A$ preserves the eigenspace with the largest eigenvalue $\|U\|$ and $A^{2}=\operatorname{Id}$ on this space. By iteration over decreasing eigenvalues of $U$ we find $A^{2}=\operatorname{Id}_{N}$. Then $A=A^{-1}=A^{*}$ has eigenvalues $\pm 1$, and if $UA>0$ there cannot be any negative eigenvalues, thus $A=\operatorname{Id}_{N}$.
Now (1.7) gives that $\partial_{t}^{j}F(0,0)=0$ for $j<m$ and $m!U(0,0)U^{*}(0,0)=\partial_{t}^{m}F(0,0)>0$, so the conditions are necessary. If $\partial_{t}^{j}F(0,0)\neq 0$ for $j<m$ then by subtracting suitable terms we obtain the following result.
Corollary 1.3.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ satisfying
(1.8) $\partial_{t}^{m}F(0,0)>0$
Then for any smooth symmetric $N\times N$ system $\mathbb{R}(t,x)$ such that $\partial_{t}^{j}\mathbb{R}(0,0)=\partial_{t}^{j}F(0,0)$, $j<m$, and $\partial_{t}^{m}\mathbb{R}(0,0)<\partial_{t}^{j}F(0,0)$, there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and ${\mathbb{M}}(t,x)$ such that $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, and
(1.9) $F(t,x)=U(t,x)\left(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x)\right)U^{*}(t,x)+\mathbb{R}(t,x)\qquad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.9) with analytic $U(t,x)$, ${\mathbb{M}}(t,x)$ and $\mathbb{R}(t,x)$. One can choose $U>0$ and then the preparation is unique.
Note that the uniqueness does not hold in general, since Remark 1.2 gives counterexamples for nonsymmetric $U$. Since (1.7) is not linear in $U(t,x)$ the proof approaches of Weierstrass or Malgrange preparation results do not work in the symmetric case, instead we shall use inverse function theorems. Theorem 1.1 will be proved in Section 3 for the analytic case, and in Section 4 for the smooth case. But first we shall consider the formal power series case.
2. Formal power series solutions
We shall assume that $F(t,x)$ is given by a formal power series expansion:
(2.1) $F(t,x)=\sum_{{j,k}}F_{j,\alpha}t^{j}x^{k}$
where $F_{j,0}=\partial_{t}^{j}F(0,0)/j!=0$ for $j<m$ and $F_{m,0}=\partial_{t}^{m}F(0,0)/m!>0$. We shall take the formal power series expansions $U(t,x)=\sum_{{j,k}}U_{\begin{subarray}{c}j,k\end{subarray}}t^{j}x^{k}$ where $|U_{0,0}|\neq 0$ and
(2.2) ${\mathbb{M}}(t,x)=\sum_{j<m}t^{j}{\mathbb{M}}_{j}(x)=\sum_{\begin{subarray}{c}j<m\\ k>0\end{subarray}}M_{j,n}t^{j}x^{k}$
is symmetric such that ${\mathbb{M}}_{j}(0)=M_{j,0}=0$, $j<m$.
Then (1.7) gives
(2.3) $\sum_{j,k}F_{j,k}t^{j}x^{k}=\sum_{j,k}U_{j,k}t^{j}x^{k}\Big(t^{m}\operatorname{Id}_{N}+\sum_{\begin{subarray}{c}j,k\end{subarray}}M_{j,k}t^{j}x^{k}\Big)\sum_{{j,k}}U_{j,k}^{*}t^{j}x^{k}$
and by matching coefficients we obtain that
(2.4) $F_{j,k}=\sum_{i,\ell}U_{i,\ell}U_{j-m-i,k-\ell}^{*}+\sum_{\begin{subarray}{c}i,\ell,p,q\end{subarray}}U_{i,\ell}M_{p,q}U_{j-p-i,k-q-\ell}^{*}$
where we sum over the nonnegative index. The first terms of this expansion are $F_{0,0}=U_{0,0}M_{0,0}U_{0,0}^{*}=0$, $F_{0,1}=U_{0,0}M_{0,1}U_{0,0}^{*}$, $F_{j,0}=0$, $j<m$, and $F_{m,0}=U_{0,0}U_{0,0}^{*}$. Here we can choose $U_{0,0}=(F_{m,0})^{1/2}$ which is the unique positive definite symmetric solution. In fact, since $U_{0,0}$ is invertible we can write any other positive symmetric solution as $U_{0,0}A$ and then $(U_{0,0}A)^{*}=A^{*}U_{0,0}$. Since $(U_{0,0}A)^{2}=U_{0,0}^{2}$, $A$ has to be orthogonal. But then Remark 1.2 gives that $A=\operatorname{Id}_{N}$. Thus, we get a unique $U_{0,0}>0$.
When $j<m$ we obtain from (2.4) that
(2.5) $U_{0,0}M_{j,k}U_{0,0}^{*}=F_{j,k}-\sum_{\begin{subarray}{c}i+\ell<j+k\end{subarray}}U_{i,\ell}M_{p,q}U_{j-p-i,k-q-\ell}^{*}$
since $|U_{0,0}|\neq 0$ we obtain symmetric $M_{j,k}$ that is given by $F_{j,k}$, $U_{i,\ell}$ with $i\leq j$ and $\ell\leq k$, and by $M_{p,q}$ with $p\leq j$, $q\leq k$ and $p+q<j+k$, Thus $M_{j,k}$ is determined by $F_{j,k}$, $U_{i,\ell}$ with $\ell<k$ or $\ell=k$ and $i<j$ and by $M_{p,q}$ with $q<k$ or $q=k$ and $p<j$.
With $j$ replaced by $j+m$, we also obtain from (2.4) that
(2.6) $2\operatorname{Re}U_{j,k}U_{0,0}^{*}=F_{j+m,k}-\sum_{\begin{subarray}{c}0<i+\ell<j+k\end{subarray}}U_{i,\ell}U_{j-i,k-\ell}^{*}-\sum_{q>0}U_{i,\ell}M_{p,q}U_{j+m-p-i,k-q-\ell}^{*}$
where the right hand side is given by $F_{j+m,k}$, $U_{i,\ell}$ with $i\leq j+m$ and $\ell<k$ or with $i+\ell<j+k$, thus by $\ell<k$ or $\ell=k$ and $i<j$, and $M_{p,q}$ with $p<m$ and $q\leq k$. Since $U_{0,0}$ is invertible, we can solve (2.6). In the case $U_{0,0}>0$, Lemma A.1 gives a unique symmetric $U_{j,k}$.
We can now determine $U_{j,k}$ and $M_{j,k}$ from (2.5) and (2.6) by induction over $j\geq 0$ for fixed $k\geq 0$ starting with $U_{0,k}$ and $M_{0,k}$ which we first determine by induction over $k$. Since $M_{0,0}=0$ we start with $U_{0,0}$ and $M_{0,1}=U_{0,0}^{-1}F_{0,1}(U_{0,0}^{*})^{-1}$, and assume that $F_{j,k}$ is known. We assume that $U_{0,\ell}$ and $M_{0,\ell+1}$ is known for $\ell<k$. Then $U_{0,k}$ is determined by these and then $M_{0,k+1}$ is determined by $U_{0,\ell}$ and $M_{0,\ell}$ for $\ell\leq k$. By induction we obtain $U_{0,k}$ and $M_{0,k+1}$ for any $k\geq 0$.
Then we use induction over $j\geq 0$ for $U_{j,k}$ and $M_{j,k+1}$ starting with $k=0$. We assume that we have determined $U_{i,\ell}$ and $M_{i,\ell+1}$ when $\ell<k$ and when $\ell=k$ for $i<j$. Then $U_{j,k}$ is determined by these and we then obtain $M_{j,k+1}$ from $U_{i,\ell}$ when $\ell\leq k$ and from $M_{i,\ell+1}$ when $\ell<k$ and when $\ell=k$ for $i<j$. By induction we obtain $U_{j,k}$ and $M_{j,k+1}$ for any $j$ and $k$. Observe that any finite expansion of $U(t,x)$ is positive definite if $|t|+|x|\ll 1$.
@@ 2. Formal power series solutions
If $F(t,x)$ is analytic (or formal power series) and symmetric such that $\partial_{t}^{m}F(0,0)>0$ and $\partial_{t}^{j}F(0,0)=0$, $\forall\,j<m$, then there exists formal power series $U(t,x)$ and ${\mathbb{M}}(t,x)$ given by (2.2) solving (1.7) modulo $O(|t|^{N}+|x|^{N})$, $\forall\,N$. We get a unique symmetric power series solution in the case when $U_{0,0}>0$.
@@
But is it not clear that in the analytic case, the formal power expansion converges to an analytic function. Instead we will use the inverse function theorem, see Section 3.
@@
It is clear that $N\times N$ analytic systems form a Banach space with the usual $L^{\infty}$ norm, and they can be localized to neighborhoods. We shall use the Taylor expansion notation
$T_{k}(F)(t,x)=\sum_{j=0}^{k-1}t^{j}\partial_{j}^{j}F(0,x)/j!$
and also the $k$ times integrations $I_{k}(F)(t,x)$ in $t$ such that $I_{1}(F)(t,x)=\int_{0}^{t}F(s,x)\,ds$ and $I_{k+1}(F)(t,x)=I_{1}(I_{k}(F))(t,x)$. Thus we find $\partial_{t}^{j}I_{k}(F)(0,x)\equiv 0$ when $j<k$ and $\partial_{t}^{k}I_{k}(F)(t,x)\equiv F(t,x)$. Observe that $t^{-k}I_{k}(F)(t,x)$ is analytic or smooth if $F(t,x)$ is.
We shall prove that if $U(t,x)$ and ${\mathbb{M}}(t,x)$ are analytic such that $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)/j!$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, then the mapping
(3.1) $(U(t,x),{\mathbb{M}}(t,x))\overset{\mathcal{F}}{\longrightarrow}\\ \Big(\operatorname{Re}\partial_{t}^{m}\big(U(t,x)(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x))U^{*}(t,x)\big),T_{m}(U{\mathbb{M}}U^{*})(t,x))\Big)\\ =(F_{1}(t,x),F_{0}(t,x))$
@@
is analytic and surjective in some neighborhood of $(0,0)$ for analytic and symmetric $F_{j}$, j=0, 1j=0,\,1, satisfying $|F_{1}(0,0)|\neq 0$, $\partial_{t}^{m}F_{0}\equiv 0$ and $F_{0}(t,0)\equiv 0$. If $\mathcal{F}(U,{\mathbb{M}})=(F_{1},F_{0})$ by (3.1) in a neighborhood of $(0,0)$ then we obtain (1.7) with
(3.2) $F(t,x)=F_{0}(t,x)+I_{m}(F_{1})(t,x)\qquad\text{near $(0,0)$}$
since $F_{0}(t,x)=T_{m}F(t,x)$ and $F_{1}(t,x)=\partial_{t}^{m}F(t,x)$.
@@
When ${\mathbb{M}}\equiv 0$ we find that $F_{0}\equiv 0$ and $\partial_{t}^{m}(t^{m}UU^{*})=m!UU^{*}=F_{1}$ when $t=0$ so $U(t,x)$ must be invertible close to $(0,0)$. For any such $U(t,x)$, we find that the differential of $\mathcal{F}$ at $(U,0)$ is given by
(3.3) $({\bf u}(t,x),{{\bf m}}(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(2t^{m}\operatorname{Re}U(t,x){\bf u}^{*}(t,x)+U(t,x){{\bf m}}(t,x)U^{*}(t,x)\big),T_{m}(U{\bf m}U^{*})(t,x)\Big)\\ =(A_{1}(t,x),A_{0}(t,x))$
where ${\bf u}(t,x)$ and ${\bf m}(t,x)$ are analytic such that $\bf m$} is symmetric, $\partial_{t}^{m}{\bf m}\equiv 0$ and ${\bf m}(t,0)\equiv 0$, which gives analytic symmetric $A_{j}$, j=0, 1j=0,\,1, such that $\partial_{t}^{m}A_{0}\equiv 0$ and $A_{0}(t,0)\equiv 0$.